OpenStudy (anonymous):

Solve (1/2)+(1/3)=1/x Explain how to do this please.

mathslover (mathslover):

oh k so first of all what is : 1/2 + 1/3 = ?

OpenStudy (anonymous):

5/6

mathslover (mathslover):

right so given : \[\large{\frac{5}{6}=\frac{1}{x}}\]

mathslover (mathslover):

cross multiply we get \[\large{x=\frac{6}{5}}\]

OpenStudy (anonymous):

Didn't think it was that simple... thanks!

mathslover (mathslover):

:P that happens no problem